The probability of a student owning a cellphone is 0.43, of owning MP 3 player is 0.38, and of owning both is 0.28. If a student is selected at random, find the probability that he or she owns a cellphone or an MP3 player.
Answer by natolino_2017(77)

let the events A:having a cellphone
B:having a MP3
P (A)= 0.43
P (B)= 0.38
P (A and B) = 0.28
P (A OR B) = P(A) + P (B) - P (A and B) = 0.43+0.38-0.28= 0.53
@natolino_